The test scores for a very large statistics class have a bell-shaped distribution with a mean of 70 points. If 16% of all students in the class scored above 85, what is the standard deviation of the scores? If 95% of the scores are between 60 and 80, what is the standard deviation?
1
Expert's answer
2021-10-25T15:23:52-0400

"\mu=70 \\

P(X>85) = 0.16 \\

1-P(X<85) = 0.16 \\

P(X<85) = 1 -0.16 = 0.84 \\

P(Z< \frac{85-70}{\sigma}) =0.84 \\

Z=0.994 \\

\frac{15}{\sigma}=0.994 \\

\sigma= 15.091"

The standard deviation of the scores is 15.091

"P(60<X<80) = 0.95 \\

P(\frac{60-70}{\sigma}<Z< \frac{80-70}{\sigma}) = 0.95 \\

P( \frac{-10}{\sigma}<Z< \frac{10}{\sigma}) = 0.95 \\

1 -2P(Z < \frac{-10}{\sigma}) = 0.95 \\

P(Z< \frac{-10}{\sigma}) = 0.025 \\

Z= -1.96 \\

\frac{-10}{\sigma} = -1.96 \\

\sigma = 5.102"

The standard deviation is 5.102